The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
10 / 60 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

10 / 60 = (1 × 10) / (6 × 10)
=1 / 6

​Thanks